# Inga Strümke5, Vince I. Madai6,7, Tobias Budig8, Ali Sunyaev9, and Adam Hilbert10 On behalf of the VALIDATE consortium

### v1

@@ V Results
Figure 4: Explainability techniques. Shades of green represent interpretable models; shades of blue represent explanation methods.
@@
The proportion of interpretability and explanation methods is shown in Fig. 4 and described in Tab. II. The most frequently used type of interpretability of FL models is algorithmic transparency, with some use of decomposability and simulatability [18]. While algorithmic transparency allows users to algorithmically trace a model’s processes, decomposability describes models of which fractions can be understood by humans. Simulatability refers to models that are sufficiently interpretable for a human to understand, or “simulate”, as a whole [18]. Among the explanation methods applied with FL models, feature relevance is predominant, followed by local explanations, and a smaller amount of model simplification-based methods [18]. Feature relevance methods quantify the impact of the model’s input features, local explanations explain specific model predictions, and simplification refers to rebuilding the entire model for easier explainability.
Agreement yes no - SUM
Definition
yes 10 7 0 17
no 7 4 9 20
SUM 17 11 9 37
TABLE III: Nomenclature and definitions.
Table III shows the amount of reviewed papers where a definition of XAI is provided and for which of these there is an agreement between their notion of XAI and the one expressed in Sec. I. Out of the 17 papers where a definition is given, 10 agree with ours and 7 do not. When a definition is not given, in 9 cases it was not possible to infer what notion of XAI is used, 7 managed the term in accordance to our definition and 4 in disagreement.
Figure 5: Amount of data used for each FL setup.
Fig. 5 shows the distribution of papers using different amounts of data for the experiments, while the colors indicate whether the FL setup was real, simulated or not specified. No papers in this survey reported real FL setups using more than 1,000,000 data points. Out of the 7 FL setups using between 100,000 and 1,000,000 data points, 3 did so in a real FL setup [73, 70, 74].
Figure 6: Number of data centers in the FL networks for each setup.
Fig. 6 shows the distribution of papers using different amounts of FL centers, while the colors indicate whether the FL setup was real, simulated, or not specified. The highest number of nodes (1,000) was simulated by [52] using standard benchmark data (MNIST, CIFAR10), followed by [9] with 100 nodes, using text datasets for language modeling; [53] with 100 nodes, using physical activity and census data from the UCI ML repository; and [75] with 66 nodes, using data from a cyberattack classification task. A minority of 9 out of 37 studies used real setups, where [76] used 12 nodes and all other papers implementing a real FL setup used 10 or less nodes. 11 papers used between 3 and 10 nodes, 2 of which using a real setup; and 9 papers used 3 or fewer centers, 5 of which using a real setup.
Figure 7: FL libraries.
Only 27.1% of the papers included in this survey used established FL libraries, whereas 37,8% developed their own libraries, and 35.1% did not specify how the network was implemented. Fig. 7 shows the FL libraries used by the analyzed papers.
Figure 8: Type of FL used.
@@
Fig. 8 illustrates the distribution of FL types used in the reviewed papers. The most prominent is horizontal FL, with 30 studies using this type exclusively. The vertical FL type is used far less, appearing in only four studies, and transfer FL is the least used, with only one study focusing on it. The overlaps in the diagram show that few studies used multiple of these methods. Only one study uses horizontal and vertical FL, another uses horizontal and transfer FL, another uses vertical and transfer FL, and one uses all three types.
Figure 9: type of data used in experiments.
Figure 10: Modality of data.
Regarding the type and modality of data, Fig. 9 shows that a majority (30%) of works experimented with real clinical data followed by a 17.5% using real finance data. 15% used widespread datasets typically oriented to model benchmarking, and the remaining works used a diversity of real data types. According to Fig. 10, the predominant modality is tabular data (52.8%) followed by images (19.4%) and time series (22.2%).
Figure 11: Number of papers that reported influence of FL on XAI compared to the amount of data used.
@@
Fig. 11 illustrates the number of papers that reported any influence between FL and XAI in relation to the amount of data used in the experiments. None of the papers using less than 1,000 data points reported any influence, whereas 9 papers working with more than 1,000 data points reported an influence.
VI Discussion
In our sample of papers jointly investigating FL and XAI,we found mostly post-hoc explanation methods. The application of explanation methods on FL systems is predominantly based on feature importance followed by local explanations, reflecting the prevalence of such methods in the literature. None of the reviewed works have proposed an ad-hoc FL method tailored to the ulterior application of explanation methods, highlighting the need for future approaches in that domain. Explanation methods designed for federated trained ML models were limited to feature aggregation [41], control of information sharing [42], and counterfactual explanations in VFL [77]. The works dealing with FL of interpretable models focused mostly on algorithmic transparency, and in the case of [60] the proposed method learns a set of interpretable rules that reflect the structure of the FL network. Interpretable models trained with FL methods were limited to fuzzy rule-based systems [60], time-series classifiers [71], SVM [78], Cox proportional hazards [79], sparse Bayesian models [80], and decision trees [69]. Fewer works deal with decomposability or simulatability, probably due to an increased difficulty in imposing those properties in an FL system.
Regarding novel methods to train interpretable ML models via FL, one very promising approach allows building federated classification models without relying on gradient descent-based methods [67]. This differs from most methods in the current literature, which focus on differentiable models. The approach of [67] is based on adapting previously existing algorithms in the boosting family to FL. Other approaches proposed an optimization method to solve the sparse SVM problem in FL [78], a novel technique called Vanishing Boosted Weights [81], or an FL network based on gradient-boosting decision trees [75].
A preference for HFL is observed in the studies surveyed, while combinations of FL types and the use of TFL are relatively rare. This is in accordance with the observed higher prevalence of HFL in the literature compared to VFL [82, 83]. Moreover, most papers apply cross-silo FL with a small set of data centers. Whether this is due to data access or this is an accurate representation of FL networks remains unclear. In general, experiments on the effect of FL on interpretability were very sparse. [60] observed an influence of FL on model interpretability in the FRBS trained using their own method, namely a larger number of model rules. Similarly, in [59], by interpreting the attention weights produced by transformer models based on a transformer architecture [84], different insights for each client were observed.
@@ VI Discussion
The fact that a majority of the reviewed papers focus on healthcare, finance, and engineering applications such as networking, highlights a high degree of interdisciplinarity and suggests cross-fertilization among fields. Conversely, this contrasts with the relative lack of studies addressing high-stakes applications in social networks, language models, and supply-chain management, where user privacy and transparent decision-making are also crucial. The need to define explicability and privacy requirements usually originates from the end users’ perspective, where data interoperability and reasoning behind automated decisions are key. The implementation of enabling technologies originates research towards using XAI for more technical goals such as improving model accuracy, assessing the training process, and prevent malicious behavior, as observed in Sec. IV-B.
In several reviewed cases, aggregating interpretability metrics from different nodes led to more general but less interpretable global insights. This fact supports the idea that FL promotes generalizability at the cost of interpretable insights from the local nodes of the FL framework. More specifically, aggregating feature relevance or attention weights across nodes can yield more generalized global insights but often sacrifices some degree of localized interpretability.
This was reported in the case of an attention mechanism [59] where each client trained locally on system log data and generated local attention weights. The global saliency map resulting from aggregating the local attention weights from the different nodes, provided a global saliency map where the contribution of individual client models is diluted, possibly leading to unique local patterns getting lost during global aggregation. The FRBS applied by [60] in a federated context involves rule generation at each data silo and merging them at the central server. Experiments observed an increased number of rules in the global model compared to local ones, indicating a growth in model complexity upon aggregating the local models, yielding a more expressive global model that is less interpretable than individual local models.
This result suggests that while FL can improve overall model performance and integrate information from diverse data sources, outputs from the global FL model could become more difficult to interpret than those from local models. This may negatively affect applications where the transparency of the local models is beneficial. One example is medical prognosis, where each local model is associated with a clinical site or hospital, and the distributions of disease features vary across hospitals. This issue could benefit from new aggregation strategies or federated explanation techniques that retain node-specific insights without compromising global model integrity.
Despite the prevalence of SHAP among the included studies, we did not find any mention of a potential federated implementation of SHAP when the supporting dataset is distributed among several data centers. Such a contribution would help exploit the representativeness of supporting data from diverse centers, potentially helping generate more accurate Shapley values. Among the studies using small amounts of data, very few reported an influence of FL on XAI, suggesting that an insufficient amount of data makes these observations challenging. It should be noted that popular feature attribution methods such as SHAP require a set of “supporting” data to do the approximate computations. Ideally, such data should not be part of the training dataset, so an insufficient number of data points probably does not allow generation of explanations robust enough to draw conclusions.
@@
Most studies were done with simulated FL networks, and the ones done with a real FL setup used datasets with less than one million samples. With ever-increasing models trained on billions of data points, the practical implications of big data remain unexplored in the context of FL and XAI. Similarly, a minority of the reviewed works use established FL libraries. The remaining papers either develop their own libraries or lack descriptions of how the network was implemented. The absence of standardized reporting for used libraries can lead to misleading observations if there are flaws in the libraries that are not identified during the analysis. A lack of transparency can hinder the development of standardized solutions, as inconsistencies in the reporting and usage of libraries complicate replicability. We also observed that many articles could improve their way of reporting data characteristics since strict adherence to reporting guidelines such as TRIPOD+AI [85] or MINIMAR [86] would ensure reproducibility and transparency. This can help standardize FL implementations, potentially accelerating the adoption of best practices in FL and XAI.
@@
The scarcity of studies explicitly addressing and quantifying the influence of FL on explanations does not give grounds for concluding that there is not such an influence, revealing an interesting research gap. Principled experimental work aimed at objectively evaluating the impact of FL training on ML model structure is also expected to shed light on its consequences on model explanations and interpretability.

### v2

@@ IV Results
Figure 3: Distribution of reviewed articles across different XAI techniques. Explanation methods (shades of blue) were studied by 64.8% of the included works, whereas interpretable models (shades of green) were studied by 35.1% of the included works.
@@
The collected interpretability and explanation methods are described in Table II. We see a slight tendency to the use of explanation methods over interpretability methods in our sample, as shown in Fig. 3. Among the explanation methods applied with FL models, feature relevance is predominant, followed by local explanations, and a smaller amount of model simplification-based methods [19]. Feature relevance methods quantify the impact of the model’s input features, local explanations explain specific model predictions, and simplification refers to rebuilding the entire model for easier explainability. In the sample of selected articles, the most frequently used type of interpretability of FL models is algorithmic transparency, with some use of decomposability and simulatability [19]. While algorithmic transparency allows users to algorithmically trace a model’s processes, decomposability describes models of which fractions can be understood by humans. Simulatability refers to models that are sufficiently interpretable for a human to understand, or “simulate”, as a whole [19].
Figure 4: Venn diagram showing the number of studies using each type of FL. Vertical and transfer FL are used far less than horizontal FL.
Definition of XAI XAI Nomenclature Count
Provided Agrees with our protocol 10
Does not agree with our protocol 7
Not provided Agrees with our protocol 7
Does not agree with our protocol 4
Unclear 9
TABLE III: Definition of explainability and nomenclature agreement
Regarding the consistency of XAI nomenclature in the literature, we observed that only 17 out of 37 articles (49.46%) provide a definition of XAI, out of which 10 agreed with our notion and 7 did not. When a definition was not given, 7 articles managed the term in accordance with our definition, 4 articles were in disagreement, and in 9 cases, it was not possible to infer which notion of XAI was used. Table III summarizes these findings.
@@
The most prominent type of FL used is horizontal, with 30 studies (81.1%) using this type exclusively. VFL is used far less, appearing in only four articles, and TFL is the least used, with only one study focusing on it. The overlaps in the diagram show that not many articles used multiple of these methods. Only one article used horizontal and transfer FL, another used vertical and transfer FL. Fig. 4 illustrates the distribution of FL types used in the reviewed studies.
Figure 5: Distribution of reviewed studies according to whether a FL library or an own implementation is used. Less than a third of the reviewed papers used established papers, and about a third did not specify how they implemented FL.
Only 10 (27.1%) articles included in this survey used established FL libraries, whereas 14 (37.8%) developed their own libraries, and 13 (35.1%) did not specify how the network was implemented. Fig. 5 shows the FL libraries used by the analyzed articles.
@@
Figure 6: Number of articles reporting and not reporting the influence of FL on XAI, grouped as a function of the amount of data used. None of the papers using less than 1,000 data points reported any influence, whereas 9 papers working with more than 1,000 data points reported an influence.
We examined the relation between the number of data points used in experiments and the reported influence between FL and XAI. Remarkably, no studies using less than 1,000 data points reported any influence, whereas 9 studies working with more than 1,000 data points reported an influence. Fig. 6 details this relation.
Figure 7: Number of studies using real, simulated, or not specified FL setups, grouped as a function of the number of data centers in the FL network. Most cases using real FL setups used 3 data centers or fewer.
Figure 8: Number of studies using real, simulated, or not specified FL setups, grouped as a function of the number of data points. No papers in this survey reported real FL setups using more than 1,000,000 data points.
No articles in this survey reported real FL setups using more than 1,000,000 data points. Out of the 7 FL setups using between 100,000 and 1,000,000 data points, 3 used a real FL setup [58, 55, 59]. Fig. 8 shows the distribution of articles using different amounts of data for the experiments, while the shading indicates whether the FL setup was real, simulated, or not specified.
The highest number of nodes (1,000) was simulated by [48] using standard benchmark data (MNIST, CIFAR10), followed by [9] with 100 nodes, using text datasets for language modeling; [60] with 100 nodes, using physical activity and census data from the UCI ML repository; and [61] with 66 nodes, using data from a cyberattack classification task. A minority of 9 out of 37 articles (24.3%) used real setups, where [62] used 12 nodes and all other studies implementing a real FL setup used 10 or fewer nodes. Eleven studies used between 3 and 10 nodes, 2 of which used a real setup, whereas 9 studies used 3 or fewer centers, 5 of which used a real setup. Fig. 7 shows the distribution of articles using different amounts of FL centers, while the colors indicate whether the FL setup was real, simulated, or not specified.
Figure 9: Distribution of reviewed studies across the different modalities of data. Tabular and time-series data are predominant, images are used in about a fifth of the studies, and a minority focused on text data.
Figure 10: Distribution of reviewed articles across different fields of application. Medicine and Life Sciences were predominant, followed by Finance and Cyber-security
Fig. 10 shows the proportion of articles across the different fields of application, where medicine and life sciences are predominant (40.5%), followed by finance (16.2%), cybersecurity (13.5%), telecommunication (8.1%) and optimization in other engineering fields (8.1%) such as electricity, mechanics, and transportation. A small proportion (13.5%) of the analyzed studies focused on the theoretical side of the ML techniques without mentioning any specific application. According to Fig. 9, the predominant modality is tabular data (51.3%) followed by images (18.1%) and time series (21.6%).
IV-A Interplay between FL and XAI
This section explores studies that have either integrated explainability into FL methods or analyzed how one technology influences the other. We first discuss instances where FL has impacted post-hoc explanations (Sec. IV-A1). Next, we examine studies where explanation methods have influenced FL training as a design step in the FL process (Sec. IV-A2). Following, we cover studies where FL has altered model interpretability (Sec. IV-A3). Finally, we discuss a study where the implementation of interpretable models has influenced FL training (Sec. IV-A4).
IV-A1 FL Impacting Explanations
Among all articles reviewed, only [43] quantitatively analyzed the impact of FL on explanations. Additionally, [63] highlights privacy concerns associated with an explanation method used on FL models.
FL is applied to taxi travel-time prediction in [43], using time series and geographical data while maintaining data privacy. The study focuses on predicting taxi trip durations in the Brunswick region with a federated deep learning model, showing that FL can achieve accuracy similar to a centralized model when synchronization is optimized. To reduce communication overhead, the authors propose a method that lowers the frequency of model updates without significantly affecting performance. They evaluate explainability in FL using attribution methods such as DeepLIFT [64], Saliency [65], Input X Gradient [66], Guided Backpropagation [67], Deconvolution [68], and Layer-wise Relevance Propagation [69]. Their results show that while local models in FL produce different variable importance rankings, aggregated feature attributions remain consistent with centralized models. This suggests that FL can incorporate explainability methods without compromising data privacy.
An FL model is trained in [63] to predict the latency in the creation of a network slice in a mobile communications scenario. Each slice manager provides data regarding CPU/RAM capacity and usage, serving as an FL node. The models are evaluated both locally and globally using SHAP [21], LIME [70], Partial Dependence Plots (PDP) [71, Ch. 8.1], and RuleFit [72]. For the PDP technique, the FL clients only share plots that display the percentage of feature impact on the target label to preserve data privacy. The influence is analyzed qualitatively with a focus on privacy concerns. Overall, it is shown that PDP explanations raise privacy concerns since they are executed on the client side.
IV-A2 Explanation Method Impacting FL Training
Among the articles that reported that the use of an explanation method impacts the FL training, a salient subset uses XAI to defend the FL process from the negative effects of defaulting nodes [35], malicious behavior from certain FL nodes [48, 60], and instances of the GAN attack in FL [73]. Other positive effects reported include improved accuracy in [45, 38, 44] and enhanced learning efficiency [74]. It must be noted that such benefits stem from incorporating XAI as a component of the FL algorithm design.
To detect malicious attacks on FL operations, [60] uses a random forest (RF) to identify features causing incorrect predictions. Each client trains both a DL and an RF model on their training data. For samples that are misclassified, it calculates the feature importance of each feature regarding incorrect classification from all decision trees and using LIME [70]. The change in a feature’s importance over time is used to assign the contribution of each feature in misclassifying the data. Such an importance value provides insights into the level of influence of each log key in a sequence during an attack, thereby indicating which features should be most protected.
A novel FL protocol was proposed in [38], where a subset of all centers available online was selected to participate in each FL round, using Shapley values (computed using SHAP) as a heuristic to estimate FL contributions from each client. More specifically, at each round, the difference between local (feature-wise) and global aggregated Shapley values at each node is used to select participating clients in an FL process. The article reports that the proposed method improves both the efficiency and accuracy compared to the FedAvg protocol [10], which does not consider client contributions.
The idea in [48] is to use explanation methods (specifically Backpropagation, Guided Backpropagation, DeepLIFT, GradCAM [42], and Integrated Gradients [75]) to detect whether each client is using malicious data to enact a backdoor attack. To this end, so-called detection filters are developed. These consist of a classifier and an explanation method that respectively identify a likely backdoor attack and triggering features in the input data. The effectiveness of various explanation methods with different classifiers is tested, strengthening the FL process against backdoor attacks. Upon testing the detection accuracy in the presence of variable proportions of backdoor attacks, the proposed methodology proves the usefulness of the different explanation methods for backdoor attack prevention.
Adaptive sparse deep networks are implemented in [74], where parameters are shared via a multi-level federated network. At each round, weights are shared at the “top” and “second” sharing levels of the FL architecture, depending on the relevance values of the network calculated through Layerwise Relevance Propagation (LRP) [76]. This approach provides good diagnostic results even when the FL dataset exhibits a non-independent and non-identically distributed (non-IID) structure.
IV-A3 FL Impacting Model Interpretability
In this section, we discuss the impact that FL training has on the interpretability of the resulting models. The featured studies, [77] and [52], discuss such an impact and its associated trade-offs.
An aggregation of client-based attention weights is investigated in [77] for a threat detection task in a cloud scenario. Using system logs as input data, each client predicts cyberattacks and computes local attention weights, which are claimed to enhance interpretability. The central server subsequently aggregates these attention weights to build a saliency map that provides insights into the impact of the different log keys on threat prediction. The influence of FL on model interpretability is assessed by comparing attention-based insights across three levels: individual cyberattacks at each client, individual attacks in the aggregated models, and aggregated attacks in the aggregated model. The aggregated insights proved to be more general but at the expense of reduced interpretability.
In [52], a fuzzy rule-based system (FRBS) [78] is trained in federation via a one-shot communication scheme. Each data silo computes its own FRBS, and the individual models are merged by the central server. The proposed FRBS uses a maximum-matching inference rule so that the inferred regression function is piecewise linear, which is inherently explainable. It is reported that FL impacts interpretability because the average number of model rules is higher when trained in the federated setting (vs. centralized). Following the intuitive idea that higher complexity implies lower interpretability, this suggests that training via FL caused lower interpretability.
IV-A4 Interpretability Impacting FL Training
Only one of the reviewed articles discussed the effects of employing interpretable models on the process of training models in federation. In the FRBS presented by [52], the inherent interpretability of the employed Takagi-Sugeno-Kang FRBS [79] models allows the central server to identify conflicts between rules inferred by data centers. This approach facilitates the reconciliation of such discrepancies, thereby enhancing the FL process.
V Discussion
@@
A majority of the reviewed studies focus on healthcare, finance, and engineering applications such as networking. This contrasts the lack of studies in other high-stakes applications, such as social networks, language models, and supply-chain management, where user privacy and transparent decision-making are also crucial. The need to define explicability and privacy requirements usually originates from the end users’ perspective, where data interoperability and reasoning behind automated decisions are key. The implementation of enabling technologies originates research toward using XAI for more technical goals, such as improving model accuracy, assessing the training process, and preventing malicious behavior, as observed in Sec. IV-A2.
A preference for HFL is evident among the studies surveyed, while combinations of FL types and applications of TFL remain relatively uncommon. This aligns with the observed higher prevalence of HFL in the literature compared to VFL [80, 81]. Additionally, most studies implement cross-silo FL involving a small set of data centers. Whether this is due to limited data access or this is an accurate representation of FL networks remains unclear.
Most studies used simulated FL networks, and the ones done with a real FL setup used datasets with less than one million data samples. This raises questions of the cause-effect relation: while data sparsity has been named as a key driver for FL [82], increasingly better-performing models are trained on billions of data points.
@@
Also surprisingly, only a minority of the reviewed studies use established FL libraries. The remaining studies either develop their own libraries or lack descriptions of how the network was implemented. The absence of standardized reporting for used libraries can cause misleading observations if there are flaws in the libraries that are not identified during the analysis. A lack of transparency can hinder the development of standardized solutions, as inconsistencies in the reporting and usage of libraries complicate replicability. We also observed that many articles could improve their way of reporting data characteristics. Strict adherence to reporting guidelines such as TRIPOD+AI [83] or MINIMAR [84] would improve reproducibility and transparency. This can help standardize FL implementations, potentially accelerating the adoption of best practices in FL and XAI.
@@
Interpretable Models Trainable in FL Explanation Methods Affected by FL Training XAI Techniques Integrated into FL Schemes
$\bullet$ Fuzzy rule-based systems [52] $\bullet$ Sparse SVMs [85] $\bullet$ Gradient Boosted Trees [61] $\bullet$ Cox regression models [86] $\bullet$ Rule-based models via boosting [87] $\bullet$ Feature relevance via SHAP, LRP [43, 44] $\bullet$ Saliency-based heatmaps (e.g., GradCAM) [48] $\bullet$ Attention weights aggregated across clients [77] $\bullet$ Local explanation differences (central vs. FL) [43] $\bullet$ Privacy concerns in client-side explanations [63] $\bullet$ Adversary detection by local explanation filters [48] $\bullet$ FL client selection via SHAP-based heuristics [38] $\bullet$ Personalized FL guided by explanation scores [44] $\bullet$ XAI-informed retraining of model components [44] $\bullet$ Feature-level privacy-preserving explanations [63]
TABLE IV: Overview of main forms of FL-XAI interplay
The remainder of this section discusses the implications of our findings regarding the interplay between FL and XAI in depth. A comprehensive summary is shown in Table IV.
V-A FL and Explanation Methods
Most of the reviewed articles employed post-hoc explanation methods, with a predominant focus on feature importance and local explanations. None of the reviewed articles have proposed an ad-hoc FL method tailored to the ulterior application of explanation methods, highlighting the need for future approaches in that domain. Explanation methods designed for federated-trained ML models were limited to feature aggregation [43], control of information sharing [63], and counterfactual explanations in VFL [88].
Despite the prevalence of SHAP among the included studies, we did not find any mention of a potential federated implementation of SHAP when the supporting dataset is distributed among several data centers. Such a contribution would help exploit the representativeness of diverse data centers, potentially helping generate more accurate Shapley values. Among the studies using small amounts of data, very few used this easy-to-measure explanation, suggesting that an insufficient amount of data makes these observations challenging. Supporting data should not be part of the training dataset, so an insufficient number of data points likely hampers the generation of robust explanations, making it difficult to draw reliable conclusions.
An important conceptual aspect that has received little attention in the reviewed literature is the distinction between local and global explanations in the context of FL. XAI refers to local explanations for individual instances and global explanations about model behavior as a whole. Global explanations may help assess whether FL results in a different model structure compared to centralized learning, potentially revealing divergences or biases in FL models. Conversely, local explanations, whether applied to client models or the aggregated model, may highlight instance-specific behaviors that differ due to local data distributions. Both levels of explanation may offer complementary insights, potentially revealing different aspects of the model’s behavior. For instance, a global explanation of a local model could help a client understand the model’s overall decision boundaries, while a local explanation of an aggregated model could clarify a prediction in a specific instance. Furthermore, structurally different models may yield similar global and local explanations if they implement equivalent decision boundaries, a phenomenon noted in neural networks due to overparameterization and symmetries [89]. Thus, whether FL leads to divergent or convergent explanations compared to centralized learning is an empirical question that remains vastly underexplored.
@@ V-A FL and Explanation Methods
The scarcity of studies explicitly addressing and quantifying the influence of FL on explanations does not give grounds for concluding that there is no such influence, revealing an interesting research gap. Principled experimental work that objectively evaluates the impact of FL training on ML model structure is expected to clarify its implications for model explanations and interpretability. Future research should also aim to clarify under which conditions FL models align or deviate from their centralized counterparts in terms of both global and local explanations and how these explanations can guide model deployment in federated contexts.
V-B FL and Interpretation Methods
The studies dealing with FL of interpretable models focused on algorithmic transparency. One proposed method learns a set of interpretable rules that reflect the structure of the FL network [52]. Fewer works deal with decomposability or simulatability, probably due to an increased difficulty in imposing those properties in an FL system. Interpretable models trained via FL methods were limited to fuzzy rule-based systems [52], time series classifiers [56], SVM [85], Cox proportional hazards [86], sparse Bayesian models [90], and decision trees [54]. A particularly promising approach enables the training of federated classification models without relying on gradient descent-based methods [51]. This differs from most established methods, which focus on differentiable models. The approach of [51] is based on adapting previously existing algorithms in the boosting family to FL. Other approaches proposed an optimization method to solve the sparse SVM problem in FL [85], a novel technique called Vanishing Boosted Weights [87], or an FL network based on gradient-boosting decision trees [61].
In general, experiments on the effect of FL on interpretability were very sparse. We identified only two articles observing an influence of FL on model interpretability [52, 77].
In [77] it was reported that aggregating interpretability metrics from different nodes led to more general but less interpretable global insights. This fact supports the idea that FL produces better generalizing models at the cost of missing some interpretable insights from the local nodes of the FL framework. Also, aggregating feature relevance or attention weights across nodes can yield more generalized global insights but can also result in sacrificing some degree of localized interpretability. This was reported in the case of an attention mechanism [77] where each client trained a local model on system log data and generated local attention weights. The global saliency map resulting from aggregating the local attention weights from the different nodes, provided a global saliency map. The contribution of individual client models in such a map was diluted, possibly leading to unique local patterns getting lost during global aggregation. The FRBS applied by [52] in a federated context involved rule generation at each data silo and merging them at the central server. Experiments observed an increased number of rules in the global model compared to local ones, indicating a growth in model complexity upon aggregating the local models, yielding a more expressive global model that is less interpretable than individual local models. These results suggest that, while FL can improve overall model performance and integrate information from diverse data sources, outputs from the global FL model could become more difficult to interpret than those from local models (trained using data from single sites). This may negatively affect applications where the transparency of the local models is beneficial. One example is medical prognosis, where each local model is associated with a clinical site or hospital, and the distributions of disease features vary across hospitals. This issue could benefit from new aggregation strategies or federated explanation techniques that retain node-specific insights without compromising global model integrity.